Let the electric equipotential lines (curves of constant potential) between two concentric cylinders with the z-axis in space be given by u(x,y) = x^2:+:y^2 = c (these are circular cylinders in xyz-space). Using the method in the text, find their orthogonal trajectories (the curves of electric force).


1
Expert's answer
2021-12-21T17:56:41-0500

Solution:


I) Differentiate the equation "x^2+y^2=c" with respect to x:



"2x+2yy^{'}=0\iff y^{'}=-\frac{x}{y}"

II) Since the equation of orthogonal trajectories is:



"y^{'}=-\frac{1}{f(x,y)}"

and in this case "f(x,y)=-\frac{x}{y}," we obtain:


"y^{'}=-\frac{1}{\frac{x}{y}}=\frac{y}{x}"

III) Now, let`s solve the differential equation:


"y^{'}=\frac{y}{x}\iff \frac{dy}{dx}=\frac{y}{x}\iff\frac{dy}{y}=\frac{dx}{x}"



IV) Integrate the left side in relation to y, and the right side in relation to x:



"\int \frac{dy}{y}= \int \frac{dx}{x} \iff ln|y|=ln|x|+c"

V) By taking exponents, we obtain:



"y=Cx"